The optical fiber has a ratio of core to cladding is 1.02 and refractive index of core is 1.50. Determine the i) critical angle, ii) Numerical aperture, iii) acceptance angle and iv) fractional index change.
1
Expert's answer
2020-04-16T09:00:21-0400

As per the given question,

The ratio in the refractive index of the core to cladding =1.02

The refractive index of the core is "(\mu_1)=1.5"

Let the refractive index of the cladding is "(\mu_2)"

hence,

"\dfrac{\mu_1}{\mu_2}=1.02"

"\mu_2=\dfrac{\mu_1}{1.02}=\dfrac{1.5}{1.02}=1.47"

a) Critical angle "C=\sin^{-1}({ \frac{\mu2}{\mu_1})}"

"C=\sin^{-1}(\frac{1.47}{1.5})=\sin^{-1}(0.98)=78.52^\circ"

b) Numerical aperture "(NA)=\sqrt{\mu_1^2-\mu^2}=\sqrt{1.5^2-1.47^2}"

"=\sqrt{0.0891}"

"=0.2984"

c)Acceptance angle "(\theta_o)=\sin^{-1}(NA)"

"=\sin^{-1}(0.2984)"

"=17.36^\circ"

d) Fractional index change

"\Delta=\dfrac{\mu_1-\mu_1}{\mu_1}"

"\Delta=\dfrac{1.5-1.47}{1.5}=0.02"